Calculate the pH of a solution made by mixing 14.5mL of 0.15M HCl with 5.3mL of 0.10M NaOH. Enter your answer to two places after the decimal.
Answer #1

Answer: -  pH = 1.08

----------------------------------------------------

Given: -

Volume of HCl = 14.5 ml = 0.0145 L

Molarity of HCl = 0.15 M = 0.15 mol/L

Volume of NaOH = 5.3 ml = 0.0053 L

Molarity of NaOH = 0.10 M = 0.10 mol/L

The reaction balanced reaction between HCl and NaOH is

  HCI + NaOH + NaCl + H2O ............(1)

Now calculate moles of HCl and NaOH

Formula: -

  Moles = Molarity x Volume(L)

Therefore,

Moles of HCl = 0.15(mol/L) 0.0145(L) = 0.002175(mol)

and

Moles of NaOH =   0.10(mol/L) x 0.0053(L) = 0.00053(mol)

Now,

We construct an ICE table for reaction (1)  

HCl + NaOH \large \rightarrow NaCl + H2O
initial 0.002175 0.00053 0 -
change - 0.00053 - 0.00053 + 0.00053 -
equilibrium 0.001645 0 0.00053 -

Total volume of Mixture = (14.5 ml + 5.3 ml) = 19.8 ml = 0.0198 L

From ICE table we have,

the moles of HCl left = 0.001645 mol

Since, HCl is a strong acid. It dissociates completely.

Thus, moles of HCl = moles of H+ ion = 0.001645 mol

Now, calculate the concentration of H+ ion in the mixture

Thus,

  = MolesOfH+Ion VolumeOf Mixture

i.e.

   0.001645(mol) (H+) = 2 = 0.0831(mol/L) = 0.0831(M) 0.0198(L)

We known that,

  pH = -log[H+

therefore,

pH = -log(0.0831) = -(-1.08) = 1.08

Therefore,

The pH of a solution made by mixing 14.5 ml of 0.15 M HCl with 5.3 ml of 0.10 M NaOH is 1.08.